# Fast LISA likelihood approximations by downsampling

### v1

@@ 7 Acknowledgements
The author is grateful to Graham Woan and John Veitch for many helpful discussions, and to Michael Williams, for help in operating his sampler Nessai. This work was partly funded by STFC grant ST/S505390/1. We are grateful for computational resources provided by Cardiff University, and funded by an STFC grant supporting UK Involvement in the Operation of Advanced LIGO. Software: dolfen is implemented in python and uses NumPy [28] and SciPy [29].
Appendix A Other required preliminary results
A.1 Noise covariance matrix introduction
The noise power generated by the detector as a function of frequency (the PSD) determines the degree to which any two samples in the time domain representation of the noise are correlated with each other. In the current section and the next, we derive the time-domain inner product operator (the matrix $\mathbf{C}^{-1}$) from the PSD to account for this interdependence of the samples. It will be assumed that the noise is wide sense stationary (WSS), that is, that its mean and covariance are time invariant.
A cross-covariance matrix of two vectors describes the covariance between elements of the vectors. The $i,j$ element of the cross-covariance matrix $\mathbf{C}_{\mathbf{XY}}$ is the covariance between the $i^{\mathrm{th}}$ element of $\mathbf{X}$ and the $j^{\mathrm{th}}$ element of $\mathbf{Y}$. A cross-covariance matrix is called an auto-covariance, or simply covariance matrix when describing the covariance between different elements of the same vector. Our noise covariance matrix will be denoted in shorthand here as $\mathbf{C}$, so that
$\mathbf{C}\equiv\mathbf{C}_{\mathbf{nn}}\,,$ (51)
where $\mathbf{n}$ is a noise realisation. In terms of matrix elements, we define a cross-covariance matrix as
$C_{ij}\equiv\mathrm{cov}[X_{i},Y_{j}]=\mathbb{E}[(X_{i}-\mathbb{E}[X_{i}])(Y_{j}-\mathbb{E}[Y_{j}])]\,,$ (52)
where $\mathbb{E}$ is the expectation value operator. For vanishing mean of the noise, we have
$C_{ij}=\mathbb{E}[n_{i}n_{j}]\equiv R_{ij}\,,$ (53)
defining the autocorrelation matrix, $R_{ij}$. Now since the noise is assumed to be WSS, the expected value of the product of any two samples only depends on the time ‘lag’ between the samples, i.e., $\mathbb{E}[n_{i}n_{j}]\!=\!\mathbb{E}[n_{i+c}n_{j+c}]$ for any integer $c$. It will be useful to consider $c=-i$ where we have that
$R_{ij}=\mathbb{E}[n_{i-i}n_{j-i}]=\mathbb{E}[n_{0}n_{j-i}]\,.$ (54)
In this form, we can see the matrix is both symmetric and circulant. It is clear that the $i,j$ element represents the expected value of noise correlations of time series samples separated by the absolute time lag between the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ sample. The first row, $R_{0i}=\mathbb{E}[n_{0}n_{i}]$, gives precisely the definition of the (discrete) ACF.
By the Wiener-Khinchin theorem, the ACF is the Fourier Transform of the PSD. However, we must be careful in the GW data analysis world since, again, we most often use the one-sided PSD which introduces a factor of 2, such that the equation embodying the Wiener-Khinchin theorem reads
$S_{\mathrm{n}}(f)=2\int_{-\infty}^{\infty}\mathrm{d}t\,R(t)e^{-2\pi ift}=2\,\mathcal{F}\{R\}(f)\,.$ (55)
Thus given the PSD we can easily compute $\mathbf{C}$, since it is defined by $R_{ij}=R(t_{i}-t_{j})$. For the inner product however, we require $\mathbf{C}^{-1}$. When $N\!=\!\dim(\mathbf{d})$ becomes large, numerical inversion of the resulting $N\times N$ noise covariance matrix becomes very unwieldy. We have devised a fast and accurate solution to this problem, which we now describe.
A.2 Fast inversion of very large autocorrelation matrices
We wish to find the matrix $\mathbf{C}^{-1}$ that satisfies $\mathbf{C}\mathbf{C}^{-1}\!=\!\mathds{1}$, where $\mathbf{C}$ is the noise covariance matrix described in the previous section. Since $\mathbf{C}$ is circulant, it follows that $\mathbf{C}^{-1}$ is circulant also. Then, in components, we can transform indices thus
$\mathds{1}_{ik}=C_{ij}C^{-1}_{jk}\longrightarrow\mathds{1}_{i^{\prime}0}=C_{i^{\prime}j^{\prime}}C^{-1}_{j^{\prime}0}=R_{i^{\prime}j^{\prime}}R^{-1}_{j^{\prime}0}\,,$ (56)
where summation over $j$ is implied by the repeated indices, and where we defined $i^{\prime}=(i-k)_{\mod N}$ and $j^{\prime}=(j-k)_{\mod N}$. Dropping primes and using the circulant property of $R_{ij}$:
$\mathds{1}_{i0}=R_{ij}R^{-1}_{j0}=R_{k0}R^{-1}_{j0}\,,$ (57)
where now $k=(i-j)_{\mod N}$. It is clear that the problem is reduced to a set of sums over only the first column from each matrix. For these column vectors, we shall use the notation $R[i]\equiv R_{i0}$, and $R^{-1}[i]\equiv R^{-1}_{i0}$. Explicitly writing the sum over $j$ from (57):
$\mathds{1}_{i0}=\sum_{j=0}^{N-1}R[(i-j)_{\mod N}]R^{-1}[j]\,.$ (58)
The vector, or discrete function $R$, is the ACF, known by the Wiener-Khinchin theorem, and where $R^{-1}$ is the (discrete) function we seek, which we call the inverse autocorrelation function.
This sum is precisely the form of a circular convolution modulo $N$ (*$*_{N}$ operator) of $R$ with $R^{-1}$, so we can write
$\mathds{1}_{i0}=(R*_{N}R^{-1})[i]\,.$ (59)
If $\mathcal{F}$ denotes the Fourier Transform and $\mathcal{F}^{-1}$ its inverse, then the convolution theorem states that
$\mathcal{F}\{R*_{N}R^{-1}\}=\mathcal{F}\{R\}\cdot\mathcal{F}\{R^{-1}\}\,,$ (60)
so that after Fourier transforming (59), using the convolution theorem, and some rearranging, we get
$R^{-1}[j]$ $=\mathcal{F}^{-1}\left\{\bigg{(}\mathcal{F}\{R\}\bigg{)}^{-1}\right\}[j]=\mathcal{F}^{-1}\left\{\left(\tfrac{1}{2}S_{\mathrm{n}}\right)^{-1}\right\}[j]\,,$ (61)
where we substituted for the PSD from (55). The discrete $R^{-1}$ is the first row of $\mathbf{C}^{-1}$, which defines the whole matrix $\mathbf{C}^{-1}$, since we know it is circulant.
A.3 Whitening transformation matrix
We will require an operator to transform vectors of correlated datapoints to vectors of uncorrelated datapoints. Suppose a noise weighted (time domain) inner product of signals can be represented in another basis in which the datapoints are uncorrelated from each other. The inner product matrix in such a basis is diagonal. Further, suppose the basis vectors are scaled such that this matrix is the identity matrix. Since inner products are invariant under basis transformations, we can define a transformed inner product of vectors $\mathbf{a}$ and $\mathbf{b}$ by
$\mathbf{a}^{\mathrm{\scriptscriptstyle T}}\mathbf{C}^{-1}\mathbf{b}=\mathbf{\bar{a}}^{\mathrm{\scriptscriptstyle T}}\mathds{1}\;\mathbf{\bar{b}}\,,$ (62)
for general time domain inner product matrix $\mathbf{C}^{-1}$, and where $\mathbf{\bar{a}}$ and $\mathbf{\bar{b}}$ are the ‘whitened’ representations of vectors $\mathbf{a}$ and $\mathbf{b}$. We shall solve this here by writing
$\mathbf{a}^{\mathrm{\scriptscriptstyle T}}\mathbf{C}^{-1}\mathbf{b}$ $=\mathbf{a}^{\mathrm{\scriptscriptstyle T}}\mathbf{C}^{(-1/2)}\mathbf{C}^{(-1/2)}\mathbf{b}$
$=\mathbf{a}^{\mathrm{\scriptscriptstyle T}}\left(\mathbf{C}^{(-1/2)}\right)^{\mathrm{\scriptscriptstyle T}}\mathds{1}\mathbf{C}^{(-1/2)}\mathbf{b}\,,$ (63)
where $\mathbf{C}^{(-1/2)}$ is some matrix to be found, which we shall constrain to be symmetric circulant (since $\mathbf{C}^{(-1)}$ is symmetric circulant, and a product of circulant matrices is circulant). We can then read off the whitening transformation of the vector space, in which $\mathbf{C}^{-1}\to\mathds{1}$ along with all vectors $\mathbf{v}\to\bar{\mathbf{v}}=\mathbf{C}^{(-1/2)}\mathbf{v}$. Noting that
$\mathbf{C}^{-1}=\mathbf{C}^{(-1/2)}\mathbf{C}^{(-1/2)}\,,$ (64)
we make use of the circulant properties to write this in components as
$C^{-1}_{ik}=C^{(-1/2)}_{ij}C^{(-1/2)}_{jk}\longrightarrow C^{-1}_{i^{\prime}0}=C^{(-1/2)}_{i^{\prime}j^{\prime}}C^{(-1/2)}_{j^{\prime}0}\,,$ (65)
where again we define $i^{\prime}=(i-k)_{\mod N}$ and $j^{\prime}=(j-k)_{\mod N}$. Dropping primes and using the circulant property of $\mathbf{C}^{(-1/2)}$:
$C^{-1}_{i0}=C^{(-1/2)}_{k0}C^{(-1/2)}_{j0}\,,$ (66)
where again we have now defined $k=(i-j)_{\mod N}$. We are now dealing with single columns of the matrices, and write these vectors as $C^{-1}[i]=C^{-1}_{i0}$, and $C^{(-1/2)}[i]=C^{(-1/2)}_{i0}$. Writing the sum in equation (66) explicitly:
$C^{-1}[i]$ $=\sum_{j=0}^{N-1}C^{(-1/2)}[(i-j)_{\mod N}]C^{(-1/2)}[j]$
$=\left(C^{(-1/2)}*_{N}C^{(-1/2)}\right)[i]\,,$
where $C^{-1}=R^{-1}$ is the inverse auto-correlation function we defined earlier in equation (61), and $C^{(-1/2)}$ shall be referred to either as the whitening function or decorrelation function and is the function sought. Fourier transforming gives
$\mathcal{F}\{C^{-1}\}=\bigg{[}\mathcal{F}\{C^{(-1/2)}\}\bigg{]}^{2}\,,$ (67)
so that
$C^{(-1/2)}=\mathcal{F}^{-1}\left\{\sqrt{\mathcal{F}\{C^{-1}\}}\right\}\,.$ (68)
However, we know that $\mathcal{F}\{C^{-1}\}=2/S_{\mathrm{n}}$, as in equation (61) for example, so we finally write the whitening function as
$C^{(-1/2)}=\mathcal{F}^{-1}\left\{\sqrt{2/S_{\mathrm{n}}}\right\}\,,$ (69)
where the matrix version $C^{(-1/2)}_{ij}\equiv C^{(-1/2)}(t_{i}-t_{j})$; this is the whitening matrix, or decorrelation matrix. From (53) and (64) we can then write
$R^{-1}_{ik}=R^{(-1/2)}_{ij}R^{(-1/2)}_{jk}$ (70)
which is the first column of, and thus defines, since it is circulant, the whitening matrix, or decorrelation matrix, $C^{(-1/2)}_{ij}=R^{(-1/2)}_{ij}$.
Appendix B Minimising the Jeffreys’ Divergence of Gaussians
@@
Let $\Theta$ be the space of parameters $\bm{\theta}$, with dimension $k=\mathrm{dim}(\Theta)$. Let $(\mathcal{D}_{\mathrm{f}},\mathbb{C}_{\mathrm{f}}^{-1})$ be the data space and inner product operator pair for the full dataset (subscript ${}_{\mathrm{f}}$). If the posterior defined on the parameter space, $p_{\mathrm{f}}(\bm{\theta})$, is Gaussian, it can be written
@@
Ideally, we would minimise the JS divergence, since this is a well-known, symmetric measure of distance between distributions and is impartial to either distribution, unlike the KL divergence, which implies its first argument is the true distribution and its second is a sample distribution. However, we are choosing our posterior to be a replacement for the truth, not some sample distribution with minimal divergence from it, so we require a symmetric divergence measure. The JS divergence is defined in equation (4.2). It contains the integral of the log of a sum and only special cases are able to be integrated analytically. In order to avoid prioritisation of any given distribution, we shall attempt to minimise another symmetric divergence known as the Jeffreys’-divergence instead, simply defined by
@@ Appendix C Numerical Marginalisation of Phase in Time-Domain
×$\hskip 20.0pt\times\mathbf{C}^{-1}\mathbf{\Big{(}d-\mathbf{A}\cos(\bm{\phi}^{\prime})\cos(\phi_{\mathrm{c}})+\mathbf{A}\sin(\bm{\phi}^{\prime})\sin(\phi_{\mathrm{c}})\Big{)}}$
@@
[…]
@@
Appendix D Downsampling visualisation with corner plots
Figure 4 shows convergence of (representatives of) posteriors as the number of time domain datapoints is increased.
Figure 4: Downsampled posteriors of the fiducial system with $\Delta f/f_{\mathrm{max}}=0.9$, $N_{\mathrm{f}}=10^{6}$ (see Section 4.4.1 and Table 4.4.2). The posteriors converge as the number of samples is increased. Selected posteriors have been included to prevent a congested plot. There is almost no discernible difference between corner plots with $N_{\mathrm{s}}\geq 256$. As $N_{\mathrm{s}}$ decreases, we can clearly see the posterior ‘breaks down’, the few samples remaining being unable to constrain the parameters in the same way as the higher $N_{\mathrm{s}}$ posteriors.

### v2

@@ 7 Acknowledgements
The author is grateful to Graham Woan and John Veitch for many helpful discussions, and to Michael Williams, for help in operating his sampler Nessai. This work was partly funded by STFC grant ST/S505390/1. We are grateful for computational resources provided by Cardiff University, and funded by an STFC grant supporting UK Involvement in the Operation of Advanced LIGO. Software: Dolfen is implemented in Python and uses NumPy [numpy] and SciPy [scipy].
Appendix A Minimising the Jeffreys’ Divergence of Gaussians
@@
Let $\Theta$ be the space of parameters $\bm{\theta}$, with dimension $k=\mathrm{dim}(\Theta)$. Let $(\mathcal{D}_{\mathrm{f}},\mathbb{C}_{\mathrm{f}}^{-1})$ be the data space and inner product operator pair for the full dataset (subscript f). If the posterior defined on the parameter space, $p_{\mathrm{f}}(\bm{\theta})$, is Gaussian, it can be written
@@
Ideally, we would minimise the Jensen-Shannon (JS) divergence, since this is a well-known, symmetric measure of distance between distributions and is impartial to either distribution, unlike the Kullback-Liebler (KL) divergence, which implies its first argument is the true distribution and its second is a sample distribution. However, we are choosing our posterior to be a replacement for the truth, not some sample distribution with minimal divergence from it, so we require a symmetric divergence measure. The JS divergence is defined in equation (4.2). It contains the integral of the log of a sum and only special cases are able to be integrated analytically. In order to avoid prioritisation of any given distribution, we shall attempt to minimise another symmetric divergence known as the Jeffreys’-divergence instead, simply defined by
@@
Appendix B Exact Preservation of Fisher Information after Downsampling
Given an $N_{\mathrm{f}}$-dimensional data space and metric pair $(\mathcal{D}_{\mathrm{f}},\mathbf{C}^{-1}_{\mathrm{f}})$ with FIM $F_{\mathrm{f},ij}$, we suppose there exists an $N_{\mathrm{s}}$-dimensional data space and metric pair $(\mathcal{D}_{\mathrm{s}},\bm{\Gamma}^{-1}_{\mathrm{s}})$ with FIM $F_{\mathrm{s},ij}$, such that
$F_{\mathrm{s},ij}=F_{\mathrm{f},ij}\,.$ (60)
Then we have
${F}_{\mathrm{s},ij}$ $={h}_{k,i}C^{-1}_{\mathrm{f},kl}{h}_{l,j}$
where primes indicate objects on the $N_{\mathrm{s}}$-dimensional data subspace, the $\bar{h^{\prime}}_{k^{\prime},i}$ are the prewhitened model derivative vectors on $\mathcal{D}_{\mathrm{s}}$, and $\bm{\Gamma}^{-1}_{\mathrm{s}}$ is the (to-be-determined) subspace metric in the whitened basis. Given a downsampling matrix $\mathbf{D}$, for example, as in equation (4), we could write
where $\omega$ is a vector to be determined, and $\mathbf{Q}$ is some guess of sample covariance matrix. Since the samples are prewhitened, we may choose $\mathbf{Q}$ to be diagonal, and take as an ansatz $\mathbf{Q}=\mathds{1}$. With this ansatz, we can set the FIM s equal (however, note there may be some other $\mathbf{Q}$ that minimises the variation between the original and downsampled posterior over the entire parameter space). Proceeding with $\mathbf{Q}=\mathds{1}$, we have that
$=\;$ $\sum_{k^{\prime}=0}^{N_{\mathrm{s}}}\omega_{k^{\prime}}^{2}\bar{h}^{\prime}_{k^{\prime},i}\bar{h}^{\prime}_{k^{\prime},j}$
$=\;$ $(\bar{h}^{\prime})^{\mathrm{\scriptscriptstyle T}}_{,i}\cdot\mathrm{diag}(\omega_{1}^{2},\omega_{2}^{2},...,\omega^{2}_{N_{\mathrm{s}}})\cdot(\bar{h}^{\prime})_{,j}\,$, (64)
which we can compare with equation (61) to see that
$\omega^{2}_{k^{\prime}}=\sum_{p=0}^{n_{b}-1}a_{p}t^{p}_{k^{\prime}}\,$, (66)
This can be written in matrix form $\mathbf{f}=\mathbf{Ba}$:
$\ell=-\tfrac{1}{2}(\mathbf{D}\bar{\mathbf{r}})^{\mathrm{\scriptscriptstyle T}}\bm{\Gamma}^{-1}_{\mathrm{s}}(\mathbf{D}\bar{\mathbf{r}})\,$, (69)
where again $\bar{\mathbf{r}}$ is the whitened residual and $\mathbf{D}$ is the sample selection matrix.
The single noise reduction factor approach discussed in the previous section is equivalent to setting the FIM s approximately equal, with the assumption that $a_{p}\!\approx\!0$ for p>0p\!>\!0 and finding the $a_{0}$ that minimises the variation in the FIM s. If a situation occurs in which this single factor is not precise enough to recover lost information, which could be expected when the Fisher information transmission rate changes significantly on timescales similar to the average duration between time samples [^1] In other words, the signal cannot be said to be slowly evolving. We can define Fisher information transmission rate as the time derivative of the expectation value of the Fisher information: $\partial_{t}\langle F_{ij}(t)\rangle$., it is possible that the solution described here will be more robust (an interesting case to consider, for example, is extreme mass ratio inspirals in LISA; the rate of transmission of Fisher information is far less uniform than slow evolution BHB s, and approximate FIM recovery using a single factor may not be feasible).
@@ Appendix C Numerical Marginalisation of Phase in Time-Domain
$\hskip 20.0pt\times\mathbf{C}^{-1}\mathbf{\Big(d-\mathbf{A}\cos(\bm{\phi}^{\prime})\cos(\phi_{\mathrm{c}})+\mathbf{A}\sin(\bm{\phi}^{\prime})\sin(\phi_{\mathrm{c}})\Big)}$
@@
$\begin{aligned} A&=\Big(\mathbf{A}\cos(\bm{\phi}^{\prime})\Big)^{\mathrm{\scriptscriptstyle T}}\mathbf{C}^{-1}\mathbf{A}\cos(\bm{\phi}^{\prime})\\ B&=-2\Big(\mathbf{A}\cos(\bm{\phi}^{\prime})\Big)^{\mathrm{\scriptscriptstyle T}}\mathbf{C}^{-1}\mathbf{\mathbf{A}\sin(\bm{\phi}^{\prime})}\\ C&=-2\mathbf{d}^{\mathrm{\scriptscriptstyle T}}\mathbf{C}^{-1}\mathbf{A}\cos(\bm{\phi}^{\prime})\end{aligned}\hskip 60.0pt\begin{aligned} D&=\mathbf{\Big(\mathbf{A}\sin(\bm{\phi}^{\prime})\Big)}^{\mathrm{\scriptscriptstyle T}}\mathbf{C}^{-1}\mathbf{A}\sin(\bm{\phi}^{\prime})\\ E&=2\mathbf{d}^{\mathrm{\scriptscriptstyle T}}\mathbf{C}^{-1}\mathbf{A}\sin(\bm{\phi}^{\prime})\\ F&=\mathbf{d}^{\mathrm{\scriptscriptstyle T}}\mathbf{C}^{-1}\mathbf{d}\,.\end{aligned}$